# Work Smarter…Not Harder! Efficient Minimization of Dependency Length in SOV Languages

### v1

@@ Introduction
Figure 1: Preverbal constituent ordering decision problem in a language with subject-object-verb (SOV) word order; Constituent (Ci) is defined as a word or a group of words that functions as a single unit within a syntactic configuration
@@
Figure 3: Dependency length and preverbal constituent ordering in Hindi (SOV); Only main verb dependencies in preverbal domain are depicted; Total dependency length (DL) and Constituent’s DL are indicated above each arc
@@ Results
In order to address the class imbalance issue for appropriate classification, we adopt an approach proposed by ? (?), originally designed for ranking web pages. Previous work dealing with the syntactic choice prediction task (?, ?), similar to our current work, has also employed this method. This transformation converts a binary classification task into a pairwise ranking task involving feature vectors of a reference sentence and each of its variants. In this case, the feature vectors consist of the input features used for each reference and variant sentence in the model. Subsequently, we train a logistic regression model on the difference between feature vectors of reference and variant pairs, as illustrated in the equations below:
@@
Table 2b presents the classification performance of our various models on held-out dataset via 10-fold cross-validation to determine how many sentences are affected by each predictor. In terms of individual performance, the model with feature ‘CL Last’, corresponding to the least-effort strategy (second column in Table 2b), accounted for the majority of our data for all SOV languages, except Basque and Japanese. In these two languages, ‘Total DL’ achieved slightly higher prediction accuracy than ‘CL Last’, with a small but statistically significant margin (p $<$ 0.001 using McNemar’s two-tailed significance test). Crucially, however, over a baseline model containing only ‘Total DL’ measure, adding the feature ‘CL Last’ (see the last column in Table 2b) induced a significant increase in classification accuracy across all SOV languages, including Basque (p $<$ 0.001 using McNemar’s two-tailed significance test compared to previous row in Table 2b). Therefore, despite the sign flip in the regression coefficient of ‘CL Last’ in the combined model for Basque (last column in Table 2a), the significant increase in 10-fold cross-validation accuracy suggests that the preverbal constituent next to the main verb, typically shorter in reference sentences than in the paired variants, continues to play an important role in Basque as well. Thus these observations highlight the effectiveness of our least-effort strategy in predicting constituent ordering choices across SOV languages.
@@ Discussion
The broader implication of our findings is that language flexibility and constituent order preferences can be understood through the lens of bounded rationality in human behavior. This behavior is constrained by cognitive abilities, available information, and the time pressure under which individuals must respond (?, ?, ?). These perspectives underscore the idea that language, as a tool for communication, has evolved to align with the cognitive abilities of its users, making it both efficient and comprehensible. Simultaneously, these perspectives also offer valuable insights into how linguistic patterns and preferences arise as a result of humans navigating these cognitive and informational limitations (?, ?).

### v2

@@ Introduction
Figure 1: Preverbal constituent ordering problem in a language with subject-object-verb (SOV) word order; Constituent (Ci) is defined as a word or a group of words that functions as a single unit within a syntactic configuration
@@
Figure 3: Preverbal constituent ordering in Hindi (SOV); Only main verb dependencies are depicted; Total dependency length (DL) and Constituent’s DL are indicated above each arc; Image reproduced from ? (?)
@@ Results
In order to address the class imbalance issue for appropriate classification, we adopt an approach proposed by ? (?), originally designed for ranking web pages. Previous work dealing with the syntactic choice prediction task (?, ?), similar to our current work, has also employed this method. This approach converts a binary classification task into a pairwise ranking task involving feature vectors of a reference sentence and each of its variants. In this case, the feature vectors consist of the input features used for each reference and variant sentence in the model. Subsequently, we train a logistic regression model on the difference between feature vectors of reference and variant pairs, as illustrated in the equations below:
@@
Table 2b presents the classification performance of our various models on held-out dataset via 10-fold cross-validation to determine how many sentences are affected by each predictor. In terms of individual performance, the model with feature ‘CL Last’, corresponding to the least-effort strategy (second column in Table 2b), accounted for the majority of our data for all SOV languages, except Basque and Japanese. In these two languages, ‘Total DL’ achieved slightly higher prediction accuracy than ‘CL Last’, with a small but statistically significant margin (p $<$ 0.001 using McNemar’s two-tailed significance test). Crucially, however, over a baseline model containing only ‘Total DL’ measure, adding the feature ‘CL Last’ (see the last column in Table 2b) induced a significant increase in classification accuracy across all SOV languages, including Basque (p $<$ 0.001 using McNemar’s two-tailed significance test compared to the previous column in Table 2b). Therefore, despite the sign flip in the regression coefficient of ‘CL Last’ in the combined model for Basque (last column in Table 2a), the significant increase in 10-fold cross-validation accuracy suggests that the preverbal constituent next to the main verb, typically shorter in reference sentences than in the paired variants, continues to play an important role in Basque as well. Thus these observations highlight the effectiveness of our least-effort strategy in predicting constituent ordering choices across SOV languages.
@@ Discussion
The broader implication of our findings is that flexibility in language and constituent order preferences can be understood through the framework of bounded rationality in human behavior. This behavior is limited by human’s cognitive abilities, information availability, and the time pressure under which individuals must respond (?, ?, ?). These perspectives emphasize that language, as a tool for communication, has evolved to align with the cognitive abilities of its users, making it both efficient and comprehensible. Simultaneously, these perspectives also offer valuable insights into how linguistic patterns and preferences emerge as a result of humans having to deal with their cognitive and informational limitations (?, ?).